OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. answer choices: 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (anonymous):

find the slope using \(m=\frac{y_2-y_1}{x_2-x_1}\) with \(x_1=1,y_1=3,x_2=4,y_2=4\)

OpenStudy (anonymous):

I need to find the equation of a line.. @satellite73

OpenStudy (anonymous):

you need the slope first

OpenStudy (anonymous):

that is, in order to find the equation of the line, you have to compute the slope as the first step then use \(y-y_1=m(x-x_1)\) and only then can you put it in the form that you want what did you get for the slope?

OpenStudy (anonymous):

m= 1/3

OpenStudy (anonymous):

ok got it !

OpenStudy (anonymous):

now use \(y-y_1=m(x-x_1)\) with \(m=\frac{1}{3}, x_1=1,y_1=3\)

OpenStudy (anonymous):

So what equation would it be?

OpenStudy (anonymous):

Umm

OpenStudy (anonymous):

I think its 3x – 3y = –8

OpenStudy (anonymous):

make the replacements i wrote and you get \[y-3=\frac{1}{3}(x-1)\]

OpenStudy (anonymous):

Or x – 3y = –8 i'm confused between the two because idk how to solve

OpenStudy (anonymous):

no it is not the one you wrote

OpenStudy (anonymous):

It's not x – 3y = –8 either? Idk what i am doing wrong.. :/

OpenStudy (anonymous):

now take \(y-3=\frac{1}{3}(x-1)\) and multiply by 3 maybe it is the second one you wrote, but it is not the first lets write it

OpenStudy (anonymous):

\[y-3=\frac{1}{3}(x-1)\] \[3y-9=x-1\] \[x-3y=-8\] yes it is that one

OpenStudy (anonymous):

Ohhh okay thanks... I have another question that is similar.. Could you help me with it? Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. answer choices: y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (anonymous):

@satellite73